A merry-go-round is initially at rest. On being given a constant angular acceleration it 

reaches an angular speed of 

1 0.50 rad/sin 10.0 s. At t = 10.0 s, determine the 

magnitude of: (i) the angular acceleration of the merry-go-round ; (ii) the linear velocity 

of a child sitting on the merry-go-round at a distance of 3.0 m from its centre; (iii) the 

tangential acceleration of the child; (iv) the centripetal acceleration of the child; and 

(v) the net acceleration of the child.


1
Expert's answer
2021-01-16T17:21:19-0500

i) By the definition of the angular acceleration, we have:


"\alpha=\dfrac{\Delta \omega}{\Delta t}=\dfrac{10.5\ \dfrac{rad}{s}}{10.0\ s}=1.05\ \dfrac{rad}{s^2}."

ii) We can find the linear velocity of the child sitting on the merry-go-round at a distance of 3.0 m from its centre from the formula:


"v=\omega r=10.5\ \dfrac{rad}{s}\cdot 3.0\ m=31.5\ \dfrac{m}{s}."

iii) We can find the tangential acceleration of the child from the formula:


"a_t=r\alpha=3.0\ m\cdot1.05\ \dfrac{rad}{s^2}=3.15\ \dfrac{m}{s^2}."

iv) We can find the centripetal acceleration of the child from the formula:


"a_c=\dfrac{v^2}{r}=\dfrac{(31.5\ \dfrac{m}{s})^2}{3.0\ m}=330.75\ \dfrac{m}{s^2}."

v) We can find the net acceleration of the child from the Pythagorean theorem:


"a_{net}=\sqrt{a_t^2+a_c^2},""a_{net}=\sqrt{(3.15\ \dfrac{m}{s^2})^2+(330.75\ \dfrac{m}{s^2})^2}=330.76\ \dfrac{m}{s^2}."

Answer:

i) "\alpha=1.05\ \dfrac{rad}{s^2}."

ii) "v=31.5\ \dfrac{m}{s}."

iii) "a_t=3.15\ \dfrac{m}{s^2}."

iv) "a_c=330.75\ \dfrac{m}{s^2}."

v) "a_{net}=330.76\ \dfrac{m}{s^2}."